# Magnitude homology of real hyperplane arrangements

### 04 Apr 2026 v1

@@ 2.3. Magnitude homology of graphs
For a graph $G$, we call a $(k+1)$-tuple $\vec{x}=(x_{0},x_{1},\ldots,x_{k})\in V^{k+1}$ of vertices of $G$ a $k$-chain of $G$ and define its length by $\ell(\vec{x})=d(x_{0},x_{1})+d(x_{1},x_{2})+\cdots+d(x_{k-1},x_{k})$. We say a $k$-chain $\vec{x}$ is proper if adjacent nodes are different $x_{0}\neq x_{1}\neq\cdots\neq x_{k}$. Let $P_{k}(G)$ be the set of proper $k$-chains of $G$, and $P_{k,\ell}(G)$ the set of proper $k$-chains of length $\ell$. The following result of Leinster is the starting point of the categorification of graph magnitude introduced by Hepworth–Willerton [16].
Proposition 2.7 (Proposition 3.9 of [25]).
For a graph $G$,
@@
In particular, $V=c_{0}$ and $E=\frac{-c_{1}}{2}$.
Let $MC_{k,\ell}(G)$ be the free abelian group generated by $P_{k,\ell}(G)$. We define the boundary operator $\partial:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ as $\partial=\sum_{i=1}^{k-1}(-1)^{i-1}\partial_{i}$, where $\partial_{i}:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ is defined by
@@
where $\widehat{x_{i}}$ means deletion of $x_{i}$. One can verify $\partial^{2}=0$. Therefore we call the homology of the chain complex $(MC_{*,\ell}(G),\partial)$ the length $\ell$ magnitude homology of $G$, denoted by $MH_{*,\ell}(G)$. Since the boundary map preserves endpoints, we have a decomposition
@@
where $MC_{k,\ell}(G;u,v)$ is the subgroup generated by chains $P_{k,\ell}(G;u,v)=\{\vec{x}\in P_{k,\ell}(G)\mid x_{0}=u,x_{k}=v\}$, forming subcomplexes and inducing decomposition of homology groups
@@
Let us denote the magnitude Betti number $\beta_{k,\ell}(G)=\mathrm{rank}MH_{k,\ell}(G)$ and the length $\ell$ magnitude Euler characteristic $\chi_{\ell}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}(G)$.
Proposition 2.8 (Propositions 2.9, 2.10 of [16]).
@@
    If $MH_{k,\ell}(G)\neq 0$, then $\ell/\mathrm{diam}(G)\leq k\leq\ell$ and the inequality $\ell/\mathrm{diam}(G)<k$ is strict if $\mathrm{diam}(G)>1$ and $\ell>0$.
The following proposition is straightforward from definition, which says magnitude can be recovered from magnitude homology.
Proposition 2.9 (Theorem 2.8 of [16]).
@@
$c_{\ell}=\chi_{\ell}(G)=\sum_{k=0}^{\infty}(-1)^{k}\beta_{k,\ell}(G)=\sum_{k=\ell/\mathrm{diam}(G)}^{\ell}(-1)^{k}\beta_{k,\ell}(G).$
@@
A direct application of the functorality is as follow. Let $G$ be a connected graph and $H$ be a gated subgraph of $G$ with gate projection $p:G\to H$. Let $\iota:H\to G$ be the inclusion. Both $p$ and $\iota$ are distance non-increasing. The composition $p\circ\iota=\text{id}_{H}$ induces $\text{id}=p_{*}\circ\iota_{*}:MH_{k,\ell}(H)\to MH_{k,\ell}(G)\to MH_{k,\ell}(H)$. This gives a decomposition of $MH_{k,\ell}(G)$.
@@
Theorem 2.12 (Theorem 5.3 of [16]).
There is a natural short exact sequence which non-naturally splits
@@
where $\square$ is the exterior product defined in Definition 5.2 of [16].
@@
For the complete graph $K_{n}$,
$MH_{k,\ell}(K_{n})=\begin{cases}\bigoplus_{\vec{x}\in P_{\ell}(K_{n})}\mathbb{Z}\vec{x},&\text{ if }k=l;\\ 0,&\text{ otherwise}.\end{cases}$
The Betti number $\beta_{\ell,\ell}(K_{n})=n(n-1)^{\ell}$. One can verify $\mathrm{Mag}(K_{n},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(K_{n})q^{\ell}=\frac{n}{1+(n-1)q}$.
@@
For the hypercube graph $Q_{d}=K_{2}^{\square d}$, we use the Künneth formula (Theorem 2.12) to obtain
@@
and use the result of $K_{2}$, we conclude
$MH_{k,\ell}(Q_{d})=\begin{cases}\mathbb{Z}^{\beta_{\ell,\ell}(Q_{d})},&\text{ if }k=l;\\ 0,&\text{ otherwise},\end{cases}$
where $\beta_{\ell,\ell}(Q_{d})=2^{d}\binom{l+d-1}{d-1}$. One can verify $\mathrm{Mag}(Q_{d},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(Q_{d})q^{\ell}=\frac{2^{d}}{(1+q)^{d}}$.
@@
We have seen that $K_{n}$ and $Q_{d}$ are diagonal. In particular the cycle graph $C_{4}=Q_{2}$ is diagonal. However Gu [15] computed $MH_{k,\ell}(C_{n})$ for all $n\geq 5$ and showed that $C_{n}(n\geq 5)$ are not diagonal.
Let us recall results from [19] in the context of graphs. Let $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$. We say $\vec{x}$ is geodesic if $\ell(\vec{x})=d(x_{0},x_{k})$. Any sub-chain of a geodesic proper chain is again geodesic proper. Any proper $1$-chain is geodesic. A proper $2$-chain $(x_{0},x_{1},x_{2})$ is geodesic if and only if $d(x_{0},x_{2})=d(x_{0},x_{1})+d(x_{1},x_{2})$.
Consider $u,v\in V(G)$, such that $\ell=d(u,v)$, then any chain in $P_{k,\ell}(G;u,v)$ must be geodesic.
@@
Let us define the length $\ell$ geodesic magnitude complex
$MC_{*,\ell}^{geod}(G):=\bigoplus_{u,v\in V(G):d(u,v)=\ell}MC_{*,\ell}(G;u,v).$
The homology is called the length $\ell$ geodesic magnitude homology, which admits the decomposition
$MH_{k,\ell}^{geod}(G)=\bigoplus_{u,v\in V(G):d(u,v)=\ell}MH_{k,\ell}(G;u,v).$
Note that $MC_{*,\ell}^{geod}(G)$ and $MH_{k,\ell}^{geod}(G)$ are trivial if $\ell>\mathrm{diam}(G)$.
Let us denote the geodesic magnitude Betti number $\beta_{k,\ell}^{geo}(G)=\mathrm{rank}MH_{k,\ell}^{geod}(G)$ and the length $\ell$ geodesic magnitude Euler characteristic $\chi_{\ell}^{geod}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}^{geod}(G)$. Notice that $\beta_{k,\ell}(G)\geq\beta_{k,\ell}^{geod}(G)$.
Kaneta–Yoshinaga [19] generalized the above geodesic magnitude chains.
Definition 2.18.
A proper $3$-chain $(x_{0},x_{1},x_{2},x_{3})\in P_{3}(G)$ is called a $4$-cut of $G$ if $(x_{0},x_{1},x_{2})$ and $(x_{1},x_{2},x_{3})$ are geodesic but $(x_{0},x_{1},x_{2},x_{3})$ is not geodesic. Write $m_{G}$ as the minimum length of $4$-cuts of $G$. We say $m_{G}=\infty$ if $G$ has no $4$-cuts.
For a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$, an internal node $x_{i}(i=1,2,\ldots,k-1)$ is called a smooth node of $\vec{x}$ if $d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$. A node that is not smooth is called a singular node. Note that endpoints are singular. We write $\varphi(\vec{x})$ as the sub-chain of $\vec{x}$ consisting of all singular nodes, called the frame of $\vec{x}$. Note that $\varphi(\vec{x})$ is not necessarily proper.
@@
Note that proper $1$-chains and proper $2$-chains are all geodesically simple, and proper chains with length smaller than $m_{G}$ are geodesically simple. Kaneta–Yoshinaga proved that the subgroups $MC_{*,\ell}^{gs}(G)$ of $MC_{*,\ell}(G)$ generated by geodesically simple chains is a subcomplex, whose homology is called the geodesically simple magnitude homology of $G$, denoted by $MH_{*,\ell}^{gs}(G)$, and the homology group $MH_{k,\ell}^{gs}(G)$ factors as a direct sum of framed magnitude homology groups as explained below.
Let $\vec{a}\in P_{m,\ell}(G)$ and $k\geq m$. We define $P_{k,\vec{a}}(G):=\{\vec{x}\in P_{k,\ell}(G)\mid\varphi(\vec{x})=\vec{a}\}$. Note that $P_{k,\vec{a}}(G)$ consists of geodesically simple chains and would be empty if $\varphi(\vec{a})\neq\vec{a}$. We further define $MC_{k,\vec{a}}(G)$ as the free abelian group generated by $P_{k,\vec{a}}(G)$, which is a subgroup of $MC_{k,\ell}^{gs}(G)$. They form a subcomplex $MC_{*,\vec{a}}(G)$ whose homology $MH_{*,\vec{a}}(G)$ is called the $\vec{a}$-framed magnitude homology of $G$. Kaneta–Yoshinaga proved the chain complex decomposes
(2.2) $MC_{k,\ell}^{gs}(G)=\bigoplus_{m\leq k,\vec{a}\in P_{m,\ell}(G)}MC_{k,\vec{a}}(G),$
and so does the homology group.
Theorem 2.20 (Theorem 3.12 of [19]).
Let $G$ be a graph and $k,\ell>0$. We have
$MH_{k,\ell}^{gs}(G)\cong\bigoplus_{m\leq k,\vec{a}\in P_{m,\ell}(G)}MH_{k,\vec{a}}(G).$
Furthermore, if $k=1$ or $\ell<m_{G}$, then $MH_{k,\ell}^{gs}(G)\cong MH_{k,\ell}(G)$.
Kaneta–Yoshinaga also related framed magnitude homology with open intervals determined by the frame. Given $a,b\in V(G)$, consider the open interval $(a,b)_{G}$ introduced in Section 2.1. Also write $C_{*}(a,b)_{G}$ the reduced chain complex of the order complex of $(a,b)_{G}$.
Proposition 2.21 (Proposition 2.3 of [13]).
For $a,b\in G$ with $d(a,b)=\ell$, there is an isomorphism of chain complexes
@@
inducing an isomorphism of homology groups
@@
For a proper chain $\vec{x}=(a,x_{1},\ldots,x_{k-1},b)\in P_{k,\ell}(G;a,b)$, it must be geodesic since $\ell=d(a,b)$. Then mapping $\vec{x}$ to the linearly order chain $x_{1}\prec_{G}\cdots\prec_{G}x_{k-1}$ in $C_{k-2}(a,b)_{G}$ defines the desired isomorphism. ∎
@@
The geodesic magnitude homology has the decomposition
$MH_{k,\ell}^{geod}(G)\cong\bigoplus_{a,b\in V(G):d(a,b)=\ell}\widetilde{H}_{k-2}(C_{*}(a,b)_{G}).$
For a frame $\vec{a}=(a_{0},a_{1},\ldots,a_{m})\in P_{m,\ell}(G)$, we denote by $C_{*}(\vec{a})_{G}$ the total complex of $C_{*}(a_{0},a_{1})_{G}\otimes\cdots\otimes C_{*}(a_{m-1},a_{m})_{G}$. A more generalized version of Proposition 2.21 is as follow.
Theorem 2.23 (Theorem 4.4 and Corollary 4.5 of [19]).
Suppose a frame $\vec{a}$ has length $\ell(\vec{a})<m_{G}$, then there is an isomorphism of chain complexes
$MC_{*,\vec{a}}(G)\to C_{*-2m}(\vec{a})_{G}$
grouping smooth nodes in order between $a_{i}$ and $a_{i+1}$, which induces isomorphisms of homology groups
$MH_{k,\vec{a}}(G)\cong\widetilde{H}_{k-2m}(C_{*}(\vec{a})_{G}).$

### 01 May 2026 v3

@@ 2.3. Magnitude homology of graphs
For a graph $G$, we call a $(k+1)$-tuple $\vec{x}=(x_{0},x_{1},\ldots,x_{k})\in V^{k+1}$ of vertices of $G$ a $k$-chain of $G$ and define its length by $\ell(\vec{x})=d(x_{0},x_{1})+d(x_{1},x_{2})+\cdots+d(x_{k-1},x_{k})$. We say that a $k$-chain $\vec{x}$ is proper if adjacent nodes are distinct, that is, if $x_{i}\neq x_{i+1}$ for all $i$. Let $P_{k}(G)$ be the set of proper $k$-chains of $G$, and let $P_{k,\ell}(G)$ be the set of proper $k$-chains of length $\ell$. The following result of Leinster is the starting point of the categorification of graph magnitude introduced by Hepworth–Willerton [17].
Proposition 2.7 ([25, Proposition 3.9]).
For a graph $G$, we have
@@
In particular, we have $c_{0}=\#V$ and $c_{1}=-2\#E$.
Let $MC_{k,\ell}(G)$ be the free abelian group generated by $P_{k,\ell}(G)$. We define the boundary operator $\partial:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ by $\partial=\sum_{i=1}^{k-1}(-1)^{i-1}\partial_{i}$, where $\partial_{i}:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ is defined by
@@
where $\widehat{x_{i}}$ denotes deletion of $x_{i}$. One can verify that $\partial^{2}=0$. We call the homology of the chain complex $(MC_{*,\ell}(G),\partial)$ the length $\ell$ magnitude homology of $G$, denoted by $MH_{*,\ell}(G)$. Since the boundary map preserves endpoints, we have a decomposition
@@
where $MC_{k,\ell}(G;u,v)$ is the subgroup generated by chains $P_{k,\ell}(G;u,v)=\{\vec{x}\in P_{k,\ell}(G)\mid x_{0}=u,\ x_{k}=v\}$. These subgroups form subcomplexes and induce a decomposition of homology groups
@@
We define the magnitude Betti number by $\beta_{k,\ell}(G)=\operatorname{rank}MH_{k,\ell}(G)$ and the length $\ell$ magnitude Euler characteristic by $\chi_{\ell}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}(G)$.
Proposition 2.8 ([17, Propositions 2.9 and 2.10]).
@@
    If $MH_{k,\ell}(G)\neq 0$, then $\ell/\operatorname{diam}(G)\leq k\leq\ell$, and the inequality $\ell/\operatorname{diam}(G)\leq k$ is strict if $\operatorname{diam}(G)>1$ and $\ell>0$.
The following proposition is a straightforward consequence of Proposition 2.7 and says that magnitude can be recovered from magnitude homology.
Proposition 2.9 ([17, Theorem 2.8]).
@@
$c_{\ell}=\chi_{\ell}(G)=\sum_{k=0}^{\infty}(-1)^{k}\beta_{k,\ell}(G)=\sum_{k=\ell/\operatorname{diam}(G)}^{\ell}(-1)^{k}\beta_{k,\ell}(G).$
@@
A direct application of functoriality is as follows. Let $G$ be a connected graph and let $H$ be a gated subgraph of $G$ with gate projection $p:G\to H$. Let $\iota:H\to G$ be the inclusion. Both $p$ and $\iota$ are distance non-increasing. The composition $p\circ\iota=\text{id}_{H}$ induces $\text{id}=p_{*}\circ\iota_{*}:MH_{k,\ell}(H)\to MH_{k,\ell}(G)\to MH_{k,\ell}(H)$. This gives a decomposition of $MH_{k,\ell}(G)$.
@@
Theorem 2.12 ([17, Theorem 5.3]).
There is a natural short exact sequence that splits non-naturally
@@
where $\square$ is the exterior product defined in [17, Definition 5.2].
@@
For the complete graph $K_{n}$, we have
$MH_{k,\ell}(K_{n})=\begin{cases}\bigoplus_{\vec{x}\in P_{\ell}(K_{n})}\mathbb{Z}\vec{x},&\text{ if }k=\ell;\\ 0,&\text{ otherwise}.\end{cases}$
The magnitude Betti number is given by $\beta_{\ell,\ell}(K_{n})=n(n-1)^{\ell}$. One can verify that $\mathrm{Mag}(K_{n},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(K_{n})q^{\ell}=\frac{n}{1+(n-1)q}$.
@@
For the hypercube graph $Q_{d}=K_{2}^{\square d}$, the Künneth formula (Theorem 2.12) gives
@@
Using the result for $K_{2}$, we obtain
$MH_{k,\ell}(Q_{d})=\begin{cases}\mathbb{Z}^{\beta_{\ell,\ell}(Q_{d})},&\text{ if }k=\ell;\\ 0,&\text{ otherwise},\end{cases}$
where $\beta_{\ell,\ell}(Q_{d})=2^{d}\binom{\ell+d-1}{d-1}$. One can verify that $\mathrm{Mag}(Q_{d},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(Q_{d})q^{\ell}=\frac{2^{d}}{(1+q)^{d}}$.
@@
We have seen in Examples 2.13 and 2.14 that $K_{n}$ and $Q_{d}$ are diagonal. In particular, the cycle graph $C_{4}=Q_{2}$ is diagonal. However, Gu [16] computed $MH_{k,\ell}(C_{n})$ for all $n\geq 5$ and showed that the graphs $C_{n}$ ($n\geq 5$) are not diagonal.
Kaneta–Yoshinaga [20] constructed a computationally tractable approximation of magnitude homology. Let $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$. We say $\vec{x}$ is geodesic if $\ell(\vec{x})=d(x_{0},x_{k})$. Every subchain of a geodesic proper chain is again geodesic and proper. Any proper $1$-chain is geodesic. A proper $2$-chain $(x_{0},x_{1},x_{2})$ is geodesic if and only if $d(x_{0},x_{2})=d(x_{0},x_{1})+d(x_{1},x_{2})$. If $u,v\in V(G)$ are vertices and $\ell=d(u,v)$, then any chain in $P_{k,\ell}(G;u,v)$ must be geodesic.
@@
We define the length $\ell$ geodesic magnitude complex by
$MC_{*,\ell}^{\mathrm{geod}}(G):=\bigoplus_{\begin{subarray}{c}u,v\in V(G)\\ d(u,v)=\ell\end{subarray}}MC_{*,\ell}(G;u,v).$
Its homology is called the length $\ell$ geodesic magnitude homology. The geodesic magnitude homology $MH^{\mathrm{geod}}_{k,\ell}(G)$ is a direct summand of $MH_{k,\ell}(G)$. It admits a decomposition
$MH_{k,\ell}^{\mathrm{geod}}(G)=\bigoplus_{\begin{subarray}{c}u,v\in V(G)\\ d(u,v)=\ell\end{subarray}}MH_{k,\ell}(G;u,v).$
Note that $MC_{*,\ell}^{\mathrm{geod}}(G)$ and $MH_{k,\ell}^{\mathrm{geod}}(G)$ are trivial if $\ell>\operatorname{diam}(G)$.
We define the geodesic magnitude Betti number by $\beta_{k,\ell}^{\mathrm{geod}}(G)=\operatorname{rank}MH_{k,\ell}^{\mathrm{geod}}(G)$ and the length $\ell$ geodesic magnitude Euler characteristic by $\chi_{\ell}^{\mathrm{geod}}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}^{\mathrm{geod}}(G)$. Observe that $\beta_{k,\ell}(G)\geq\beta_{k,\ell}^{\mathrm{geod}}(G)$.
Geodesic magnitude homology is a coarse approximation of magnitude homology. As a more refined approximation, Kaneta–Yoshinaga [20] introduced geodesically simple magnitude homology. For a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$, an internal node $x_{i}$ ($i=1,2,\ldots,k-1$) is called a smooth node of $\vec{x}$ if $d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$. A node that is not smooth is called a singular node. Note that endpoints are singular. We write $\varphi(\vec{x})$ for the subchain of $\vec{x}$ consisting of all singular nodes and call it the frame of $\vec{x}$. Note that $\varphi(\vec{x})$ is not necessarily proper.
@@
Definition 2.19.
We define $MC_{k,\ell}^{\mathrm{gs}}(G)$ to be the subgroup of $MC_{k,\ell}(G)$ generated by geodesically simple chains. Then $MC_{*,\ell}^{\mathrm{gs}}(G)$ forms a subcomplex of $MC_{k,\ell}(G)$ [20, Propositions 3.6 and 3.7]. Its homology is called the geodesically simple magnitude homology of $G$, denoted by $MH_{*,\ell}^{\mathrm{gs}}(G)$.
How well geodesically simple magnitude homology approximates magnitude homology is controlled by the notion of a $4$-cut.
Definition 2.20.
A proper $3$-chain $(x_{0},x_{1},x_{2},x_{3})\in P_{3}(G)$ is called a $4$-cut of $G$ if $(x_{0},x_{1},x_{2})$ and $(x_{1},x_{2},x_{3})$ are geodesic but $(x_{0},x_{1},x_{2},x_{3})$ is not geodesic. Let $m_{G}$ be the minimum length of $4$-cuts of $G$. We set $m_{G}=\infty$ if $G$ has no $4$-cuts. Note that $m_{G}\geq 3$. Any proper chain with length smaller than $m_{G}$ is geodesically simple.
Lemma 2.21.
If $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$ satisfies $\varphi(\vec{x})=(x_{0},x_{k})$ and $\vec{x}$ is not geodesically simple, i.e. $(x_{i-1},x_{i},x_{i+1})$ is geodesic for $i=1,\ldots,k-1$, but $\vec{x}$ is not geodesic, then $\vec{x}$ has a subchain that is a $4$-cut.
Proof.
Let $\vec{y}=(y_{0},\ldots,y_{p})$ be a contiguous subchain of $\vec{x}$, that is, $y_{0}=x_{i},y_{1}=x_{i+1},\ldots,y_{p}=x_{i+p}$ for some $i$, such that $\vec{y}$ is not geodesic and $p$ is as small as possible. This means $(y_{0},\ldots,y_{p-1})$ and $(y_{1},\ldots,y_{p})$ are geodesic. Then $p\geq 3$. We claim $\vec{z}:=(y_{0},y_{1},y_{p-1},y_{p})$ is a $4$-cut. First note that $y_{p-1}\neq y_{1}$ as they are connected by a geodesic $(y_{1},y_{2},\ldots,y_{p-1})$. That $(y_{0},y_{1},y_{p-1})$ and $(y_{1},y_{p-1},y_{p})$ are geodesic follows from the assumption that $(y_{0},\ldots,y_{p-1})$ and $(y_{1},\ldots,y_{p})$ are geodesic. Moreover, $\ell(\vec{z})=d(y_{0},y_{1})+d(y_{1},y_{p-1})+d(y_{p-1},y_{p})=\sum_{i=1}^{p}d(y_{i-1},y_{i})>d(y_{0},y_{p})$, since $\vec{y}$ is not geodesic. This shows $\vec{z}$ is not geodesic and hence is a $4$-cut. ∎
Let $\vec{a}\in P_{m,\ell}(G)$ and $k\geq m$. We define $P_{k,\vec{a}}(G):=\{\vec{x}\in P_{k,\ell}(G)\mid\varphi(\vec{x})=\vec{a}\}$. Note that $P_{k,\vec{a}}(G)$ consists of geodesically simple chains and is empty if $\varphi(\vec{a})\neq\vec{a}$. We further define $MC^{\mathrm{gs}}_{k,\vec{a}}(G)$ as the free abelian group generated by $P_{k,\vec{a}}(G)$, which is a subgroup of $MC_{k,\ell}^{\mathrm{gs}}(G)$. These groups form a subcomplex $MC^{\mathrm{gs}}_{*,\vec{a}}(G)$ whose homology $MH^{\mathrm{gs}}_{*,\vec{a}}(G)$ is called the $\vec{a}$-framed magnitude homology of $G$. Kaneta–Yoshinaga proved that the chain complex $MC_{*,\ell}^{\mathrm{gs}}(G)$ decomposes as
(2.3) $MC_{k,\ell}^{\mathrm{gs}}(G)=\bigoplus_{\begin{subarray}{c}m\leq k\\ \vec{a}\in P_{m,\ell}(G)\end{subarray}}MC^{\mathrm{gs}}_{k,\vec{a}}(G),$
and the same is true for the homology groups.
Theorem 2.22 ([20, Theorem 3.12]).
Let $G$ be a graph and $k,\ell>0$. We have a decomposition
$MH_{k,\ell}^{\mathrm{gs}}(G)\cong\bigoplus_{\begin{subarray}{c}m\leq k\\ \vec{a}\in P_{m,\ell}(G)\end{subarray}}MH^{\mathrm{gs}}_{k,\vec{a}}(G).$
Furthermore, if $k=1$ or $\ell<m_{G}$, then $MH_{k,\ell}^{\mathrm{gs}}(G)\cong MH_{k,\ell}(G)$.
Kaneta–Yoshinaga also related framed magnitude homology to open intervals determined by the frame. Given $a,b\in V(G)$, consider the open interval $(a,b)_{G}$ introduced in Section 2.1. We also write $C_{*}(a,b)_{G}$ for the reduced chain complex of the order complex of $(a,b)_{G}$.
Proposition 2.23 ([13, Proposition 2.3]).
For $a,b\in V(G)$ with $d(a,b)=\ell$, there is an isomorphism of chain complexes
@@
which induces an isomorphism of homology groups
@@
Any proper chain $\vec{x}=(a,x_{1},\ldots,x_{k-1},b)\in P_{k,\ell}(G;a,b)$ must be geodesic since $\ell=d(a,b)$. Mapping $\vec{x}$ to the linearly ordered chain $x_{1}\prec_{G}\cdots\prec_{G}x_{k-1}$ in $C_{k-2}(a,b)_{G}$ defines the desired isomorphism. ∎
@@
The geodesic magnitude homology admits a decomposition
$MH_{k,\ell}^{\mathrm{geod}}(G)\cong\bigoplus_{\begin{subarray}{c}a,b\in V(G)\\ d(a,b)=\ell\end{subarray}}\widetilde{H}_{k-2}(C_{*}(a,b)_{G}).$
For a frame $\vec{a}=(a_{0},a_{1},\ldots,a_{m})\in P_{m,\ell}(G)$, we denote by $C_{*}(\vec{a})_{G}$ the total complex of $C_{*}(a_{0},a_{1})_{G}\otimes\cdots\otimes C_{*}(a_{m-1},a_{m})_{G}$. A more general version of Proposition 2.23 is as follows.
Theorem 2.25 ([20, Theorem 4.4 and Corollary 4.5]).
Suppose a frame $\vec{a}$ has length $\ell(\vec{a})<m_{G}$. Then there is an isomorphism of chain complexes
$MC^{\mathrm{gs}}_{*,\vec{a}}(G)\to C_{*-2m}(\vec{a})_{G},$
which induces isomorphisms of homology groups
$MH^{\mathrm{gs}}_{k,\vec{a}}(G)\cong\widetilde{H}_{k-2m}(C_{*}(\vec{a})_{G}).$
Remark 2.26.
The original statement of Theorem 4.4 in [20] does not contain the assumption $\ell(\vec{a})<m_{G}$. The result fails if a frame node $a_{i}$ of $\vec{a}$ can be smoothed out by nodes from neighboring intervals. For example, let $G=C_{4}$ be the cycle graph of $4$ vertices, labeled by $0,1,2,3$ in cyclic order. Choose $\vec{a}=(0,1,3)\in P_{2,3}(G)$. Then $(0,1)_{G}$ is empty and $(1,3)_{G}$ consists of two incomparable points $\{0,2\}$. Compare $MC^{\mathrm{gs}}_{3,\vec{a}}(G)$ with $C_{-1}(\vec{a})_{G}$. The element $\varnothing\otimes 0$ in $C_{-1}(\vec{a})_{G}$ does have a preimage $(0,1,0,3)$ in $MC^{\mathrm{gs}}_{3,\vec{a}}(G)$. However, the element $\varnothing\otimes 2$ in $C_{-1}(\vec{a})_{G}$ does not have a preimage because $(0,1,2,3)$ is a $4$-cut.
Proof of Theorem 2.25.
Let $\vec{a}=(a_{0},a_{1},\ldots,a_{m})\in P_{m,\ell}(G)$ such that $\varphi(\vec{a})=\vec{a}$. Then $\ell=\ell(\vec{a})<m_{G}$. If $\vec{x}\in P_{k,\vec{a}}(G)$ is a basis element of $MC^{\mathrm{gs}}_{k,\vec{a}}(G)$, then $\vec{x}$ has the form
$\vec{x}=(a_{0},x^{1}_{1},\ldots,x^{1}_{k_{1}},a_{1},x^{2}_{1},\ldots,x^{2}_{k_{2}},a_{2},\ldots,a_{m}).$
Note that $k=m+\sum_{i=1}^{m}k_{i}$. The smooth nodes between $a_{i-1}$ and $a_{i}$ form a linearly ordered chain $\langle x^{i}\rangle:=(x^{i}_{1}\prec_{G}\cdots\prec_{G}x^{i}_{k_{i}})$ in $C_{k_{i}-1}(a_{i-1},a_{i})_{G}$. Define $\Phi_{k}:MC^{\mathrm{gs}}_{k,\vec{a}}(G)\to C_{k-2m}(\vec{a})_{G}$ by
$\Phi_{k}(\vec{x})=\langle x^{1}\rangle\otimes\langle x^{2}\rangle\otimes\cdots\otimes\langle x^{m}\rangle.$
The map $\Phi_{k}$ is injective on basis elements. To prove that $\Phi_{k}$ is also surjective, consider a basis element $\langle y^{1}\rangle\otimes\cdots\otimes\langle y^{m}\rangle$ of $C_{k-2m}(\vec{a})_{G}$, where $\langle y^{i}\rangle=(y^{i}_{1}\prec_{G}\cdots\prec_{G}y^{i}_{k_{i}})\in C_{k_{i}-1}(a_{i-1},a_{i})_{G}$ for $i=1,\ldots,m$. Let
$\vec{y}=(a_{0},y^{1}_{1},\ldots,y^{1}_{k_{1}},a_{1},y^{2}_{1},\ldots,y^{2}_{k_{2}},a_{2},\ldots,a_{m}).$
We have $\vec{y}\in P_{k,\ell}(G)$. It remains to prove $\varphi(\vec{y})=\vec{a}$. First note that $(a_{i-1},y^{i}_{1},\ldots,y^{i}_{k_{i}},a_{i})$ is geodesic (2.1), and hence $y^{i}_{j}$ is smooth in $\vec{y}$. Next suppose, for contradiction, that $a_{i}$ is smooth in $\vec{y}$ for some $i=1,\ldots,m-1$, that is,
$d(y^{i}_{k_{i}},y^{i+1}_{1})=d(y^{i}_{k_{i}},a_{i})+d(a_{i},y^{i+1}_{1}).$
Here $y^{i}_{k_{i}}$ is understood as $a_{i-1}$ if $k_{i}=0$. Similarly $y^{i+1}_{1}$ is understood as $a_{i+1}$ if $k_{i+1}=0$. Note that $k_{i}$ and $k_{i+1}$ cannot both be zero, since otherwise $\varphi(\vec{a})\neq\vec{a}$. Consider the subchain
$\vec{z}=(a_{i-1},y^{i}_{1},\ldots,y^{i}_{k_{i}},a_{i},y^{i+1}_{1},\ldots,y^{i+1}_{k_{i+1}},a_{i+1}).$
It satisfies $\varphi(\vec{z})=(a_{i-1},a_{i+1})$ and $\ell(\vec{z})=d(a_{i-1},a_{i})+d(a_{i},a_{i+1})>d(a_{i-1},a_{i+1})$ by the assumption that $a_{i}$ is singular in $\vec{a}$. By Lemma 2.21, $\vec{z}$ contains a subchain $\vec{w}$ that is a $4$-cut. However, $\ell(\vec{w})\leq\ell(\vec{z})\leq\ell(\vec{y})=\ell$, which contradicts the assumption $\ell<m_{G}$. The verification that $\Phi_{k}$ defines a chain map is omitted. ∎